OpenStudy (anonymous):

1. solve. -2/5x - 9<9/10 2. solve. 4x + 6 < -6

OpenStudy (campbell_st):

1. add 9 to both sides of the equation \[\frac{-2}{5} x < 9\frac{9}{10}\] next multiply both sides by 5 \[-2x < 5 \times (9 \frac{9}{10})\] lastly, divide by -2 and reverse the inequality for the solution.

OpenStudy (anonymous):

ok thanks what about # 2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well subtract 6 from both sides then divide by 4